OpenStudy (vishweshshrimali5):

A fair die is rolled four times. The probability that each of the final three rolls is at least as large as the roll preceeding it is m/n in the lowest terms. Then n -10m = ?

OpenStudy (anonymous):

Please stop spamming and wait patiently for a tutor to assist you with your problems.

OpenStudy (vishweshshrimali5):

KKKKKKK sorry

OpenStudy (vishweshshrimali5):

Hey!! Xcuseme Who are you polpak ? May I know please?

OpenStudy (anonymous):

Just a lego guy

OpenStudy (vishweshshrimali5):

What is lego guy ?

OpenStudy (vishweshshrimali5):

I mean what does it mean ?

OpenStudy (vishweshshrimali5):

Hey You have 580 medals !!!!!!!!!! wow

OpenStudy (anonymous):

this question is a poser

OpenStudy (anonymous):

Others have more.

OpenStudy (vishweshshrimali5):

Hello satellite please help me !

OpenStudy (vishweshshrimali5):

What is a poser !

OpenStudy (anonymous):

are u studying iit coaching

OpenStudy (vishweshshrimali5):

No but preparing for that but how do u know that man ?

OpenStudy (anonymous):

if the first roll is a1, then probability is 1 if the first roll is a 2 then probability is \[(\frac{5}{6})^3\]

OpenStudy (anonymous):

that type of question comes in iit

OpenStudy (vishweshshrimali5):

Yaa u are right amingad !

OpenStudy (anonymous):

it is easy take 4 turn as 36c4

OpenStudy (vishweshshrimali5):

How do u got that satellite please explain !

OpenStudy (anonymous):

oh wait that is wrong. i need to read more carefully

OpenStudy (vishweshshrimali5):

Please take you time

OpenStudy (vishweshshrimali5):

How do you got that amingad

OpenStudy (anonymous):

looks like coupon collectors problem

OpenStudy (vishweshshrimali5):

It may be

OpenStudy (anonymous):

what is probability of a record. gives harmonic series

OpenStudy (anonymous):

yikes now i have to think

OpenStudy (vishweshshrimali5):

Amingad please explain step by step I can't understand it so fastly

OpenStudy (vishweshshrimali5):

I am a neophyte in this field you may think that

OpenStudy (anonymous):

fair dice consist of 6 numbers . so the the no of times get the highest probability of number will be 24 c4 my mistake i wrote 36 c4

OpenStudy (anonymous):

you lost me here for sure

OpenStudy (vishweshshrimali5):

What do you mean to say satellite ?

OpenStudy (vishweshshrimali5):

Okkkkk amingad now is there any step left ?

OpenStudy (anonymous):

i mean i am lost completely.

OpenStudy (vishweshshrimali5):

Ohhhh you mean that it is a nice question then thanks a lot

OpenStudy (vishweshshrimali5):

Well satellite you can have a look at other questions

OpenStudy (anonymous):

12c3< m/ n so n has to be highest dice to be rolled

OpenStudy (vishweshshrimali5):

Amingad will you please do me a favour

OpenStudy (anonymous):

yep it is tricky but good

OpenStudy (vishweshshrimali5):

can you write the answer in a single post and with each step explained Because my teacher says that you will have to answer the question with each step explained or you will get no marks

OpenStudy (vishweshshrimali5):

Will you please do so ?

OpenStudy (vishweshshrimali5):

Are you answering the question amingad

OpenStudy (zarkon):

\[\frac{7}{72}\]

OpenStudy (anonymous):

yep correct answer zarkon

OpenStudy (zarkon):

i know

OpenStudy (zarkon):

Just for fun...if we used an n-sided die the probability is \[\frac{(n+1)(n+2)(n+3)}{24n^3}\]

OpenStudy (zarkon):

and if we rolled an n-sided die k times it would be \[\frac{n(n+1)(n+2)\cdots(n+k-1)}{n^{k}k!}\]